# Irreducibility and periodicity in $\mathbb{Z}^{2}$ symbolic systems

### v1

@@ 3.5 Application to rectangles
Then $D\in\mathcal{D}_{n}$ are $20$-sparse sets of diamonds of dimensions $4w_{n}\times 4h_{n}$. Define $\operatorname{safe}(\mathcal{R}_{1},\ldots,\mathcal{R}_{N})=\operatorname{safe}(\mathcal{D}_{1},\ldots,\mathcal{D}_{N})$. Statements (2)–(4) are now immediate consequences of the corresponding statements in Proposition 3.9.
@@ 4.1 Notation and definitions
      The long sides of $R$ are parallel to $u$, and when the long sides are extended in direction $u$, the resulting ray passes within $\frac{1}{100}r$ of the center of $D$.
@@ 4.2 Parameters
By Theorem LABEL:prop:coins-and-buckets, there exist unorientable coin-and-bucket configurations with $1000^{2}n^{2}$ buckets. We introduce two sequences of real parameters
@@ 4.3 Certificates and their parts
    The coordinates of $w\in W$ are real, but from $w$ a site in $\mathbb{Z}^{2}$ by taking the coordinate-wise integer part.
To help foordinate this data we shall introduce additional geometric structures:
@@
    • The center of an $n$-frame $F$ is denoted $c(F)$
@@
    If the set of centers of $n$-frames is $10r_{n}$-dense, we say that the $n$-th level is dense. The certificate is dense if all its levels are dense. We allow non-dense certificates, which arrise naturally in the intermediate stages of constructing a certificate.
@@
    • Every two distinct boxes $B_{i}\neq B_{j}$ satisfy $d(B,B^{\prime})\geq 100h_{n}$.
    • All the $B_{i}$ in a frame have a common orientation, and furthermore the orientations lies in the set
@@
    • Each $n$-box $B$ associated to a frame $F$ is divided into $1000$ rectangular sections $B^{(1)},B^{(2)},\ldots,B^{(1000)}$ of dimensions $(w_{n}/1000)\times h_{n}$, numbered by increasing distance from the center of $F$.
@@
      $\mathcal{R}_{C}(B)=\{B^{\prime},B^{\prime\prime}\}\cup\left\{R^{(i)}\left|\begin{array}[]{c}\text{$R$\text{ is a $k$-box for }}k<n\text{ such that }\theta(R)=\theta(B)\\ \text{and }2R^{(i)}\cap B^{(i)}\neq\emptyset\text{ for some }i\in\{1,\ldots,1000\}\end{array}\right.\right\}$
@@
Figure 3: Part of a frame $F$ and a box $B$ in $F$ (indicated in red), and two lower level frames with boxres in the same orientation (indicated in blue and yellow). The gray rectangles indicate sections of the lower-level boxes which would be included in $\mathcal{R}_{C}(B)$. The witness $w$ of $B$ is indicated by a black dot, and the red path passing through $w$ consists entirely of safe points relative to $\mathcal{R}_{C}(B)$.
  • The set of all witnesses associated to a frame $F$ denoted
@@
4.4 Comments and first corrolaries of the definition
@@
We implement this a follows. Each witness belongs to some section $S$, which has an associated position in its box ($\sigma(S)\in\Sigma$), and the box has an orientation ($\theta(S)\in\Theta$). This partitions witnesses into $|\Theta|\cdot|\Sigma|=1000^{2}$ classes parameterized by $\Theta\times\Sigma$. One should think of these classes as occupying different “layers” of the plane, avoiding each other. We must still ensure that witnesses from each class do not come close to each other. This is done in the following lemmas:
@@
If $F,F^{\prime}$ are $n$-frames whose centers are at least $r_{n}/2$ separated, and if $B\in F$, $B^{\prime}\in F^{\prime}$ are boxes with $\theta(B)=\theta(B^{\prime})$, then $d(B^{(i)},(B^{\prime})^{(i)})>\frac{1}{4}r_{n}$ for all $i\in\{1,\ldots,1000\}$
@@
Let $p,p^{\prime}$ denote the points in $B,B^{\prime}$ closest to the centers of $F,F^{\prime}$ respectively. Since the rectangles are almost radial, $d(p,c(F))<r_{n}/100$ and $d(p^{\prime},c(F^{\prime}))<r_{n}/100$. Since $d(c(F),c(F^{\prime}))>r_{n}/2$, we have $d(p,p^{\prime})>r_{n}/3$. Each section is of dimensions $(w_{n}/1000)\times h_{n}$ and $h_{n}<w_{n}=\frac{9}{10}r_{n}$, so the diamater $d$ of a section is at most $d<\sqrt{2}\cdot w_{n}<r_{n}/100$. Finally, $\theta(B)=\theta(B^{\prime})$, so for each $i\in\Sigma$ there are points $p_{i}\in B^{(i)}$ and $p^{\prime}_{i}\in(B^{\prime})^{(i)}$ such that $d(p_{i},p^{\prime}_{i})=d(p,p^{\prime})>r_{n}/3$. It follows that
@@
Let $C$ be a certificate and $w,w^{\prime}\in C$ distinct witnesses belonging to levels $n\neq n^{\prime}$, respectively. If $\sigma(w)=\sigma(w^{\prime})$ and $\theta(w)=\theta(w^{\prime})$, theSn $d(w,w^{\prime})\geq h_{\min\{n,n^{\prime}\}}$.
@@
  • $\theta(w)\neq\theta(w^{\prime})$
  • $\sigma(w)\neq\sigma(w^{\prime})$
@@
If $w,w^{\prime}$ come from different frames in the same level, then Lemma 4.1 says that either $\theta(w)\neq\theta(w^{\prime})$ or $\sigma(w)\neq\sigma(w^{\prime})$.
@@
The following lemma will be useful when we need to define boxes in a new frame, and want them to intersect a set $E$ (the set where we want to place witnesses, because we have reedom to define the symbols there).
@@
For every $n$-frame $F$ and every path-connected set $E\subseteq F\setminus$$\frac{1}{5}F$ of diameter $>\frac{1}{100}r_{n}$, there are $N_{n}+1$ rectangles $R_{1},\ldots,R_{N_{n}+1}$ such that
@@
Without loss of generality we can assume that $E$ is a simple path. Choose $u,v\in\Theta$ that differ by an angle of $10002\pi/$ 1000 and such that the projection of $E$ to $u^{\perp}$ and $v^{\perp}$ both lie within $r_{n}/50$ of the corresponding projection of $c(F)$. It is elementary to see that there is a universal constant $c>0$ such that one of these projections of $E$ is an interval of length at least $cr_{n}$. Suppose $u$ is this direction. Since $r_{n}\gg h_{n},n$ we may assume the projection of $E$ in direction $u^{\perp}$ has length is at least $1000(N_{n}+1)h_{n}$. This allows us to construct $N_{n}+1$ almost radial rectangles in direction $u$ satisfying (1)-(3) and such that their projection to $u^{\perp}$ lies inside the projection of $E$, and this ensures also (4). ∎
@@ 4.5 The alphabet
Recall that $\Theta$ is the set of angles that are multiples of $10002\pi/$ 1000 and $\Sigma=\{1,\ldots,1000\}$ is the index set of the sections of boxes. Our alphabet will be
@@ 4.6 Compatibility
Recall that we have defined witnesses to be points in $\mathbb{R}^{2}$, but would like to use them to index configurations in To this end, for $w=(w_{1},w_{2})\in\mathbb{R}^{2}$ write $[w]=([w_{1}],[w_{2}])$, where []delimited-[][$\cdot$][ ⋅ ] denotes the integer part. Then we define
@@ 4.7 Definition of the subshift
Fix $n\in\mathbb{N}$; it suffices to show that $x$ is $n$-aperiodic. Let $F$ be an $n$-frame in $C$. Then $x$ is compatible with $F$, so $CBC(x,F)$ is unorientable and, in particular, there is a bucket $(n,\theta,\sigma)$ in $CBC(x,F)$ contains coins of different orientations. If $w,w^{\prime}\in F$ are witnesses in corresponding to these coins, then the fact that they are in the same bucket means that $w=w^{\prime}\bmod n$, $\theta(w)=\theta(w^{\prime})=\theta$ and $\sigma(w)=\sigma(w^{\prime})=\sigma$. The fact that the coins have different orientations means that $\pi_{\theta,\sigma}(x_{w})\neq\pi_{\theta,\sigma}(x_{w^{\prime}})$. Thus $x_{w}\neq x_{w^{\prime}}$, and $x$ not $n$-periodic. ∎
@@
One can correct this flaw as follows: rather than defining witnesses to be points in $\mathbb{R}^{2}$, define them as pairs $(w,s)$ where $w\in\mathbb{R}^{2}$ and $s\in\mathbb{Z}^{2}$ with $\left\|w-s\right\|_{\infty}\leq 1/2$, and use $s$ in place of $[w]$ in the definition of compatibilty. With this change, the space of certificates is still compact, and convergence of witnesses includes convergence of the “integer parts” $s$, so compatibility does pass to the limit. and the proof above goes through.
@@
Fix a certificate $C$ (it is not hard to see that, taking all $h_{n},w_{n}$ large relative to their predecessors, certificates exist).
By Corollary 4.3, witnesses $w,w^{\prime}\in C$ with the same orientation and section must satisfy $d(w,w^{\prime})>w_{1}>2$ and hence the integer points derived from $w,w^{\prime}$ satisfy $[w]\neq[w^{\prime}]$.
@@ 5.1 Reduction to bounded, simply connected regions
(B) $X$ admits gluing along all pairs $E,F\subseteq\mathbb{Z}^{2}$ with $d(E,F)>g$ and $F$ finite.
[…]
Next, we claim that (B) follows from
(C) $X$ admits gluing along all all pairs $E,F\subseteq\mathbb{Z}^{2}$ such that
@@
    $F$ is finite and $E$ has finitely many connected components.
Indeed, if $E,F$ satisfy the assumptions of (B), and if $x,y\in X$, replace $E$ by $E^{\prime}=\{u\in\mathbb{Z}^{2}\mid d(u,F)>g\}$. Since $F$ is finite, $E^{\prime}$ has finitely many connected components, so by (C) there is a gluing $z$ of $x|_{E^{\prime}}$ and $y|_{F}$. Since $E\subseteq E^{\prime}$, the configuration $z$ is also a gluing of $x|_{E}$ and $y|_{F}$.
We can assume the connected components are $\frac{1}{2}g-2$ separated
Specifically, we claim that (C) follows from
@@
    $F$ is finite and $E$ has finitely many connected components.
  3. D3.
Indeed, suppose that $E,F$ satisfy the conditions (C1) and (C2). We will enlarge and merge pairs of components to obtain sets $E^{\prime},F^{\prime}$ containing $E,F$, respectively, in a way that (D1)-(D3) are satisfied; then (D) ensures gluing along $E^{\prime},F^{\prime}$, which implies gluing along $E,F$.
To this end, view $E,F$ as subsets of $\mathbb{R}^{2}$, and enlarge $E$ by adding to it all line segments between pairs $u,v\in E$ that are neighbors in $\mathbb{Z}^{2}$. Do the same for $F$. Note that subsets of $E\cap\mathbb{Z}^{2}$ that were previously connected components in the induced graph from $\mathbb{Z}^{2}$ are now connected components in the topological sense, and that $E,F$ are still $g$-separated after adding these line segments.
Given any $u,v\in E\cap\mathbb{Z}^{2}$ that lie in different connected components of $E$ and satisfy $d(u,v)<\frac{1}{2}g$, introduce a line segment $\ell=\ell_{u,v}$ connecting $u,v$. Observe that if $w\in\mathbb{Z}^{2}$ and $d(w,\ell)\leq\frac{1}{2}g$, then the point $w^{\prime}\in\ell$ closest to $w$ is within $g/2$ of one of the points $u,v$, and therefore either $d(w,u)<g$ or $d(w,v)<g$.
Let $\widehat{E}$ denote the union of $E$ and all line segments $\ell_{u,v}$ introduced in the previous paragraph to $u,v\in E$, and similarly $\widehat{F}$. If $I\subseteq\widehat{E}$ is a connected component, then for any $u\in I$ we have $d(u,w)>\frac{1}{2}g$ for $w\in\widehat{F}$ and $w\in\widehat{E}\setminus I$ : Indeed, we showed this above for $w\in(F\cup(E\setminus I))\cap\mathbb{Z}^{2}$, but it holds also for $w\in\widehat{F}\cup(\widehat{E}\setminus I)$ because the distance between a pair of line segments is attained at their endpoints, which in our case belong to $E$ or $F$).
We now return to $\mathbb{Z}^{2}$: For each line segment $\ell$ in $\widehat{E}$ add to $E$ the integer points in a $1$-neighborhood of $\ell$. In the resulting set $E^{\prime}$, each connected component (in the graph $\mathbb{Z}^{2}$) is obtained from a connected component of $\widehat{E}$, and the former lies in the $1$-neighborhood of the latter. Fo the same to obtain $F^{\prime}$ from $\widehat{F}$. Now $E\subseteq E^{\prime}$, $F\subseteq F^{\prime}$, and $E^{\prime},F^{\prime}$ satisfy (D1) and (D3); they also satisfy (D2) because $E,F$ satisfy (C2).
We can assume that $E,F$ are connected
This is because condition (D) above follows from
(E) $X$ admits gluing along pairs $E,F$ provided
  1. E1.
    $d(E,F)>\frac{1}{2}g-2$.
  2. E2.
    $F$ is finite and $E,F$ are connected.
  3. E3.
For suppose that $E,F$ satisfy the conditions (D1)-(D3) and assume that (E) holds. We show that every $x,y\in X$ admit a gluing along $E,F$. We do so by induction on the total number of connected components in $E,F$. For, assuming as we may that one of the sets $E,F$ has more than one component, let $J$ be one of these components. If $J\subseteq E$, apply the induction hypothesis to obtain a gluing $z$ of $x,y$ along $E\setminus J$ and $F$ (which now have one component less), and then glue $x$ and $z$ along $J$ and $J^{\prime}=\{u\in\mathbb{Z}^{2}\mid d(u,J)>\frac{1}{2}g-2\}$. The resulting configuration $w$ is also a gluing of $x,y$ along $E,F$. In the case $J\subseteq F$, first glue $x,y$ along $E,F\setminus J$, and glue the result and $y$ along $J^{\prime}$ and $J$
We can assume that $F$ is simply connected and $E=\{u\in\mathbb{Z}^{2}\mid d(u,F)>\frac{1}{2}g-2\}$
This reduction is not strictly necessary, but it gives something more concrete to think about. In fact, we can achieve this simply by “filling in the holes” in $F$, and then setting $E$ as above.

### v2

@@ 4.5 Application to rectangles
Then $D\in\mathcal{D}_{n}$ are $20$-sparse sets of diamonds of dimensions $4w_{n}\times 4h_{n}$. Define $\operatorname{safe}(\mathcal{R}_{1},\ldots,\mathcal{R}_{N})=\mathcal{H}(\mathcal{D}_{1},\ldots,\mathcal{D}_{N})$. Statements (2)–(4) are now immediate consequences of the corresponding statements in Proposition 4.9.
@@ 5.1 Notation and definitions
      The long sides of $R$ are have length $\leq r$, are parallel to $u$, and when they are extended in direction $u$, each of the resulting rays passes within $\frac{1}{100}r$ of the center of $D$.
@@
    Note that $\theta(u)$ points approximately from $R\cap\partial D$ to the center of $D$.
@@ 5.2 Parameters
By Proposition 3.1, there exist unorientable coin-and-bucket configurations with $1000^{2}n^{2}$ buckets. We introduce two sequences of real parameters
@@ 5.3 Certificates and their parts
    The coordinates of $w\in W$ are real, but from $w$ we obtain a site in $\mathbb{Z}^{2}$ by taking the coordinate-wise integer part.
To help coordinate this data we shall introduce additional geometric structures:
@@
    • The center of an $n$-frame $F$ is denoted $c(F)$ .
@@
    If the set of centers of $n$-frames is $10r_{n}$-dense, we say that the $n$-th level is dense. The certificate is dense if all its levels are dense. We allow non-dense certificates, which arise naturally in the intermediate stages of constructing a certificate.
@@
    • Every two distinct boxes $B_{i}\neq B_{j}$ satisfy $d(B_{i},B_{j})\geq 100h_{n}$.
    • All the $B_{i}$ in a frame have a common orientation, and furthermore the orientations lie in the set
@@
    • Each $n$-box $B$ associated to a frame $F$ is divided into $1000$ closed rectangular sections $B^{(1)},B^{(2)},\ldots,B^{(1000)}$ of dimensions $(w_{n}/1000)\times h_{n}$, numbered by increasing distance from the center of $F$.
      We require the sections to be closed because later a limiting procedure will require it. Evidently as described above the sections cannot be disjoint, but one should think of them as such. Formally one can insert minuscule spaces between them, while slightly reducing their length in the radial direction. We do not bother to keep track of this detail which presents no serious problems.
@@
[…]
@@
Figure 3: Part of a frame $F$ and a box $B$ in $F$ (indicated in red), and two lower level frames with boxes in the same orientation (indicated in blue and yellow). The gray rectangles indicate sections of the lower-level boxes which would be included in $\mathcal{R}_{C}(B)$. The witness $w$ of $B$ is indicated by a black dot, and the red path passing through $w$ consists entirely of safe points relative to $\mathcal{R}_{C}(B)$.
  • The set of all witnesses associated to a frame $F$ is denoted
@@
5.4 Comments and first corollaries of the definition
@@
We implement this as follows. Each witness belongs to some section $S$, which has an associated position in its box ($\sigma(S)\in\Sigma$), and the box has an orientation ($\theta(S)\in\Theta$). This partitions witnesses into $|\Theta|\cdot|\Sigma|=1000^{2}$ classes parameterized by $\Theta\times\Sigma$. One should think of these classes as occupying different “layers” of the plane, avoiding each other. We must still ensure that witnesses from each class do not come close to each other. This is done in the following lemmas:
@@
If $F,F^{\prime}$ are $n$-frames whose centers are at least $r_{n}/2$ separated, and if $B\in F$, $B^{\prime}\in F^{\prime}$ are boxes with $\theta(B)=\theta(B^{\prime})$, then $d(B^{(i)},(B^{\prime})^{(i)})>\frac{1}{4}r_{n}$ for all $i\in\{1,\ldots,1000\}$ .
@@
Let $p,p^{\prime}$ denote the points in $B,B^{\prime}$ closest to the centers of $F,F^{\prime}$ respectively. Since the rectangles are almost radial, $d(p,c(F))<r_{n}/100$ and $d(p^{\prime},c(F^{\prime}))<r_{n}/100$. Since $d(c(F),c(F^{\prime}))>r_{n}/2$, we have $d(p,p^{\prime})>r_{n}/3$. Each section is of dimensions $(w_{n}/1000)\times h_{n}$ and $h_{n}\ll w_{n}=\frac{9}{10}r_{n}$, so (assuming as we may that $h_{n}<w_{n}/1000$), the diameter $d$ of a section is at most $d\leq\sqrt{2}w_{n}/1000<r_{n}/100$. Finally, $\theta(B)=\theta(B^{\prime})$, so for each $i\in\Sigma$ there are points $p_{i}\in B^{(i)}$ and $p^{\prime}_{i}\in(B^{\prime})^{(i)}$ such that $d(p_{i},p^{\prime}_{i})=d(p,p^{\prime})>r_{n}/3$. It follows that
@@
Let $C$ be a certificate and $w,w^{\prime}\in C$ distinct witnesses belonging to levels $n\neq n^{\prime}$, respectively. If $\sigma(w)=\sigma(w^{\prime})$ and $\theta(w)=\theta(w^{\prime})$, then $d(w,w^{\prime})\geq h_{\min\{n,n^{\prime}\}}$.
@@
  • $\theta(w)\neq\theta(w^{\prime})$ .
  • $\sigma(w)\neq\sigma(w^{\prime})$ .
@@
If $w,w^{\prime}$ come from different frames in the same level, then either $\theta(w)\neq\theta(w^{\prime})$, or else, by Lemma 5.1, $\sigma(w)\neq\sigma(w^{\prime})$.
@@
The following lemma will be useful when we need to define boxes in a new frame, and want them to intersect a set $E$ (the set where we want to place witnesses, because we have freedom to define the symbols there).
@@
For every $n$-frame $F$ and every path-connected set $E\subseteq F\setminus$$\frac{1}{5}F$ of diameter $\frac{1}{100}r_{n}$, there are $N_{n}+1$ rectangles $R_{1},\ldots,R_{N_{n}+1}$ such that
@@
Without loss of generality we can assume that $E$ is a simple path. Choose $u,v\in\Theta$ that differ by an angle of $2\pi/1000$ and such that the projection of $E$ to $u^{\perp}$ and $v^{\perp}$ both come within $r_{n}/50$ of the corresponding projection of $c(F)$. It is elementary to see that there is a universal constant $c>0$ such that one of these projections of $E$ is an interval of length at least $cr_{n}$. Suppose $u$ is this direction. Since $r_{n}\gg h_{n},n$ we may assume the projection of $E$ in direction $u^{\perp}$ has length is at least $1000(N_{n}+1)h_{n}$. This allows us to construct $N_{n}+1$ almost radial rectangles in direction $u$ satisfying (1)-(3) and such that their projection to $u^{\perp}$ lies inside the projection of $E$, and this ensures also (4). ∎
@@ 5.5 The alphabet
Recall that $\Theta$ is the set of angles that are multiples of $2\pi/1000$ and $\Sigma=\{1,\ldots,1000\}$ is the index set of the sections of boxes. Our alphabet will be
@@ 5.6 Compatibility
Recall that we have defined witnesses to be points in $\mathbb{R}^{2}$, but would like to use them to index configurations in $\mathbb{Z}^{2}$. To this end, for $w=(w_{1},w_{2})\in\mathbb{R}^{2}$ write $w]=([w_{1}],[w_{2}])$, where $\cdot$] denotes the integer part. Then we define
@@ 5.7 Definition of the subshift
Fix $n\in\mathbb{N}$; it suffices to show that $x$ is $n$-aperiodic. Let $F$ be an $n$-frame in $C$. Then $x$ is compatible with $F$, so $CBC(x,F)$ is unorientable and, in particular, there is a bucket $(n,\theta,\sigma)$ in $CBC(x,F)$ containing coins of different orientations. If $w,w^{\prime}\in F$ are witnesses corresponding to these coins, then the fact that they are in the same bucket means that $w=w^{\prime}\bmod n$, $\theta(w)=\theta(w^{\prime})=\theta$ and $\sigma(w)=\sigma(w^{\prime})=\sigma$. The fact that the coins have different orientations means that $\pi_{\theta,\sigma}(x_{w})\neq\pi_{\theta,\sigma}(x_{w^{\prime}})$. Thus $x_{w}\neq x_{w^{\prime}}$, and $x$ not $n$-periodic. ∎
@@
One can correct this flaw as follows: rather than defining witnesses to be points in $\mathbb{R}^{2}$, define them as pairs $(w,s)$ where $w\in\mathbb{R}^{2}$ and $s\in\mathbb{Z}^{2}$ with $\left\|w-s\right\|_{\infty}\leq 1/2$, and use $s$ in place of $[w]$ in the definition of compatibility. With this change, the space of certificates is still compact, and convergence of witnesses includes convergence of the “integer parts” $s$, so compatibility does pass to the limit. and the proof above goes through.
@@
Fix a dense certificate $C$ (it is not hard to see that, taking all $h_{n},w_{n}$ large relative to their predecessors, dense certificates exist).
By Corollary 5.3, witnesses $w,w^{\prime}\in C$ with the same orientation and section must satisfy $d(w,w^{\prime})>w_{1}>2$ and hence the integer points derived from $w,w^{\prime}$ satisfy $w]\neq[w^{\prime}]$.
@@ 6.1 Reduction to bounded, simply connected regions
(B) $X$ admits gluing along all pairs $E,F\subseteq\mathbb{Z}^{2}$ with $d(E,F)>g$ and $E,F$ are finite (and in particular, have finitely many connected components).
For, given arbitrary $E,F$ with $d(E,F)\geq g$, and given $x,y\in X$, for each $n$ we can use (B) to find a gluing $z_{n}$$\in X$ of $x|_{E\cap[-n,n]^{2}}$ and $y|_{F\cap[-n,n]^{2}}$. Any accumulation point of $(z_{n})$ is then a gluing of $x|_{E}$ and $y|_{F}$.
We can assume connected components are $\frac{1}{2}g-2$ separated
Specifically, we claim that (B) follows from
(C) $X$ admits gluing along pairs $E,F$ that satisfy
@@
    $E,F$ have are finite.
  3. C3.
Indeed, suppose that $E,F$ satisfy (B). We will enlarge and merge pairs of components to obtain sets $E^{\prime},F^{\prime}$ containing $E,F$, respectively, in a way that (C1)-(C3) are satisfied; then (C) ensures gluing along $E^{\prime},F^{\prime}$, which implies gluing along $E,F$.
To this end, view $E,F$ as subsets of $\mathbb{R}^{2}$, and enlarge $E$ by adding to it all line segments between pairs $u,v\in E$ that are neighbors in $\mathbb{Z}^{2}$. Do the same for $F$. Note that subsets of $E\cap\mathbb{Z}^{2}$ that were previously connected components in the induced graph from $\mathbb{Z}^{2}$ correspond to components in the topological sense, and that $E,F$ are still $g$-separated after adding these line segments.
Given any $u,v\in E\cap\mathbb{Z}^{2}$ that lie in different connected components of $E$ and satisfy $d(u,v)<\frac{1}{2}g$, introduce a line segment $\ell=\ell_{u,v}$ connecting $u,v$. Let $\widehat{E}$ denote the union of $E$ and all line segments $\ell_{u,v}$ as above. Define $\widehat{F}$ analogously.
Claim 6.1.
Suppose that $\ell_{u,v}$ and $\ell_{u^{\prime},v^{\prime}}$ are two line segments with $d(u,v)<\frac{1}{4}g$ and $d(u^{\prime},v^{\prime})<\frac{1}{4}g$. Then
  1. 1.
    If $w\in\mathbb{R}^{d}$ and $d(\ell_{u,v},w)<\frac{1}{4}g$ then $d(\{u,v\},w)<\frac{1}{2}g$.
  2. 2.
    If $d(\ell_{u,v},\ell_{u^{\prime},v^{\prime}})<\frac{1}{4}g$ then $d(\{u,v\},\{u^{\prime},v^{\prime}\})<\frac{1}{4}g$.
Proof.
(1) Since the length of $\ell_{u,v}$ is less than $\frac{1}{2}g$, every $z\in\ell_{u,v}$ is at distance of no more than $\frac{1}{4}g$ from one of the endpoints. Therefore, if $d(w,\ell)\leq\frac{1}{4}g$, and choosing $z\in\ell_{u,v}$ to be the point nearest $w$, we have
$d(w,\{u,v\})\leq d(w,z)+d(z,\{u,v\})<\frac{1}{4}g+\frac{1}{4}g=\frac{1}{2}g$
(2) If $d(\ell_{u,v},\ell_{u^{\prime},v^{\prime}})<\frac{1}{4}g$, then the distance is attained a pair of points $z\in\ell_{u,v}$ and $z^{\prime}\in\ell_{u^{\prime},v^{\prime}}$ with either $z\in\{u,v\}$ or $z^{\prime}\in\{u^{\prime},v^{\prime}\}$. In the first case, apply (1) with $w=z$ to conclude $d(z,\{u^{\prime},v^{\prime}\})<\frac{1}{2}g$, and $z\in\{u,v\}$ gives us $d(\{u,v\},\{u^{\prime},v^{\prime}\})<\frac{1}{2}g$ as claimed. The other case follows similarly. ∎
Since $\widehat{E},\widehat{F}$ differ from $E,F$ only by the addition of segments $\ell_{u,v}$ as above between connected components, it follows from the claim that if $I\subseteq\widehat{E}$ is a connected component, then for any $u\in I$ we have $d(u,w)>\frac{1}{2}g$ for $w\in\widehat{F}\cup(\widehat{E}\setminus I)$.
We now return to $\mathbb{Z}^{2}$: For each line segment $\ell$ in $\widehat{E}$ add to $E$ the integer points in a $1$-neighborhood of $\ell$. In the resulting set $E^{\prime}$, each connected component (in the graph $\mathbb{Z}^{2}$) is obtained from a connected component of $\widehat{E}$, and the former lies in the $1$-neighborhood of the latter. Do the same to obtain $F^{\prime}$ from $\widehat{F}$. Now $E\subseteq E^{\prime}$, $F\subseteq F^{\prime}$, and $E^{\prime},F^{\prime}$ satisfy (C1) and (C3); they also satisfy (C2) because $E,F$ satisfy (C2).
We can assume that $E,F$ are connected
Condition (C) above follows from
@@
    $E,F$ are finite and connected.
For suppose that $E,F$ satisfy the conditions (C1)-(C3) and assume that (D) holds. We show that every $x,y\in X$ admit a gluing along $E,F$. We do so by induction on the total number of connected components in $E,F$. When there are two components altogether this is exactly (D). Now suppose the implication is known when there are up to $n$ components, and assume $E\cup F$ have $n+1$ components. Consider the partial order where components $J,J^{\prime}$ satisfy $J<J^{\prime}$ if $J^{\prime}$ separates $J$ from infinity. Let $J$ be a connected component of $E\cup F$ that is minimal with respect to this order and observe that all other components of $E\cup F$ are contained in the unbounded component $J^{\prime}$ of $\{u\in\mathbb{R}^{d}\mid d(u,J)>\frac{1}{2}g-2\}$
Since $d(E,F)>0$, either $J\subseteq E$ or $J\subseteq F$. Suppose for instance that $J\subseteq F$, the other case is analogous.
If $J=F$, then we simply use (D) to glue $x,y$ along $J^{\prime},F$, and this is also a gluing along $E,F$, as required.
If $J\neq F$ then $F^{\prime}=F\cap J^{\prime}$ has fewer connected components than $F$ and $E,F^{\prime}$ still satisfies (C) so by the induction hypothesis we can glue $x,y$ along $E,F^{\prime}$ to obtain $z$. Now we apply (D) to glue $z$ and $y$ along $J^{\prime},J$, yielding a gluing of $x,y$ along $E,F$.
One more (not strictly necessary) reduction
Condition (D) follows from
(E) $X$ admits gluing along pairs $E,F$ that satisfy
  1. 1.
    $F$ is finite and simply connected.
  2. 2.
    $E$ is the unbounded component of $\{u\in\mathbb{Z}^{2}\mid d(u,F)>\frac{1}{2}g-2\}$.
This reduction is not  necessary for the later proofs, but it gives a simpler situation to think about. To see that (E) implies (D), suppose that $E,F$ are as stated in (D).
If $F$ does not separate $E$ from infinity, fill in all the holes in $F$ to obtain a finite simply connected set $F^{\prime}$, let $E^{\prime}$ denote the unbounded component of $\{u\in\mathbb{Z}^{2}\mid d(u,F^{\prime})>\frac{1}{2}g-2\}$, and note that $E\subseteq E^{\prime}$. Therefore by (E) we can glue along $E^{\prime},F^{\prime}$ and this gives a gluing along $E,F$.
If $F$ does separate $E$ from infinity, swap the roles of $E,F$, and proceed as above.